OpenStudy (anonymous):

Is this true always, sometimes, or never? (Points : 1) Always Sometimes Never 10. Is this true always, sometimes, or never? x ÷ 0 = 0 (Points : 1) Always Sometimes Never 11. A program on television ran 30 minutes. The program ran seven minutes longer than three times the length of the commercials. Which equation represents this problem? (Points : 1) 3x + 7 = 30 x + 7 = 3(30) 3x = 30 + 7

OpenStudy (anonymous):

10 is never cant have division by 0

OpenStudy (anonymous):

number 9 is x*x=x to the power of two

OpenStudy (nubeer):

9 true. always.

OpenStudy (nubeer):

3x+7 =30

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thanks! how bout these? 12. A program on television ran 30 minutes. The program ran seven minutes longer than three times the length of the commercials. Use the guess and check method to find the number of minutes the commercials ran during the program. (Points : 1) 12.3 3.7 7.7 13. The Sears Tower is 1,454 feet tall. Let the height of the antenna on the tower be represented by the variable a. The total height of the building plus the antenna is 1,707 feet. Which equation represents this problem? (Points : 1) 1,454 = 1,707 + a 1,454+a=1,707 1,454 - a = 1,707 14. The Sears Tower is 1,454 feet tall. Let the height of the antenna on the tower be represented by the variable a. The total height of the building plus the antenna is 1,707 feet. What is the height of the antenna? (Points : 1) 253 feet 353 feet 450 feet 15. Write an equation to represent the problem. The area of a triangle is equal to one half times the base times the height. Use A for area, b for base and h for height. (Points : 1) A = 1/2 b + h A = 1/2 (b + h) A = (1/2) bh 16. Which equation represents the problem? The circumference of a circle equals the product of two, pi, and the radius. Use C for circumference, 3.14 for pi, and r for the radius. (Points : 1) C = 2 + 3.14 + r C = 2(3.14)r C = (3.14r)/2 17. Which equation represents the problem? A cube has six sides, each of which is a square. If the measure of each side is s, what is the area of one side (square) of the cube? (Points : 1) A = s^3 A = s^2 A = 6s^2 18. Write a formula for the entire surface area (all six squares) of a cube in simplest form. Let SA be the surface area and s is the measure of one edge of the cube. (Points : 1) SA = s^3 SA = s^2 SA = 6s^2

OpenStudy (nubeer):

17. B 18. C

OpenStudy (nubeer):

Most of the questions are simple english.. it does not involve maths in it.. i am sure u can do too if u try..

OpenStudy (anonymous):

i can but i have 15 other projects to do and turn in by 8am this morning im not gonna sleep tonight so i just need help on these please

OpenStudy (nubeer):

ok i will do it this time.. 16. B 15. C 14. A 13. B 12. C

OpenStudy (anonymous):

thank you :)

OpenStudy (nubeer):

you're welcome :)